# The Unbounded Denominators Conjecture

### v3

@@ 4.6. An enhancement of Ihara’s Lemma
(4.6.1) $H^{1}(\Gamma(N),\mathbf{F}_{q})^{\oplus 2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q}),$
@@
    The twisted restriction map coming from viewing $\psi\in H^{1}(\Gamma(N),\mathbf{F}_{q})$ as a map from $\Gamma(N)\rightarrow\mathbf{F}_{q}$ and then considering the map
    $A\psi:\Gamma(N)\cap\Gamma_{0}(p),\quad g\mapsto\psi(AgA^{-1}).$
@@
The kernel of the map (4.6.1) lies inside $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{q})^{2}$.
@@
This version of Ihara’s Lemma was essentially proved by Ribet in [Rib84]. The proof is just an abelian version of Lemma 4.4.1. We recall some of the details. Let $\Phi\subset\Gamma(N)$ be the maximal normal subgroup whose quotient is an elementary $q$-abelian group $T$. Canonically, we have $\Gamma(N)/\Phi\simeq T$ and $H^{1}(\Gamma,\mathbf{F}_{q})\simeq\mathrm{Hom}(T,\mathbf{F}_{q})$. The kernel of the map
@@
Exactly as in the proof of Lemma 4.4.1, we deduce from Goursat’s Lemma that the cokernel $\Delta$ arises from two maps from $\Gamma(N)$ to $\Delta$ which agree along $\Gamma(N)\cap\Gamma_{0}(p)$, and thus on their amalgam $\mathrm{SL}_{2}(\mathbf{Z}[1/p])(N)$. Since $\mathrm{SL}_{2}(\mathbf{Z}[1/p])(N)$ as the congruence subgroup property, it thus arises from a congruence quotient of this group at primes away from $p$. But that precisely means that the classes in $\mathrm{Hom}(T,\mathbf{F}_{q})=H^{1}(\Gamma(N),\mathbf{F}_{q})$ become trivial after passing to a congruence subgroup $\Gamma^{\prime}\subset\Gamma(N)$, hence the claim. ∎
@@
also lies inside $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{q})^{2}$.
@@ 5. The uniformization of $\mathbf{C}\smallsetminus\mu_{N}$
In this section we develop all the particular analytic properties that we need of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ for $N\geq 2$. André has pointed out to us that our two main results here, Theorem 5.1.4 and Lemma 5.2.17, appear in work of Kraus and Roth [KR16, Remark 5.1 and Theorems 1.2 and 1.10]. Nevertheless, as our proofs are simplified to cover our current needs, and since the results of Kraus and Roth rely on some previous work of themselves and others, we keep our self-contained exposition as a convenience to the reader, and refer to [ASVV10, KRS11, KR16] and the references there for various further results and a more thorough study of the uniformization of $\mathbf{C}\smallsetminus\mu_{N}$. The reader will also benefit from the material in § III.1 in Goluzin’s book [Gol69], which recovers $F_{N}$ via an explicit computation of the Riemann map of a $\mathbf{Z}/N\mathbf{Z}$-rotationally symmetric circular $N$-gon, taking the case of zero angles and doing Schwarz reflections in the sides of the circular polygon.
@@ 5.1. Schwarzians and the conformal radius
where $\widetilde{\Gamma}_{N}\subset\mathrm{PSL}_{2}(\mathbf{R})$ denotes the Fuchsian group of $\mathbf{C}\smallsetminus\mu_{N}$. The map $\widetilde{F}_{N}$ is unique up to the action of $\mathrm{PSL}_{2}(\mathbf{R})$ on the source.
We now pin down the definition of $\widetilde{F}_{N}$ with the following precise choices. Using the action of $\mathrm{PSL}_{2}(\mathbf{R})$ we may assume that $\widetilde{F}_{N}(i)=0$. The stabilizer of $i\in\mathbf{H}$ consists of Möbius transformations given by elements in $\mathrm{SO}_{2}(\mathbf{R})$, so by specifying $\widetilde{F}_{N}(i\infty)=1$ (here by abuse of notation, we also use $\widetilde{F}_{N}$ to denote the natural extension of $\widetilde{F}_{N}$ on $\mathbf{H}$ to cusps), we determine $\widetilde{F}_{N}$ uniquely.
@@
Note that $F_{N}$ is just the map $\widetilde{F}_{N}$ composed with the standard conformal isomorphism $D(0,1)\rightarrow\mathbf{H}$ sending 0 to $i$, and hence
@@
(5.1.3) $\widetilde{r}_{N}=\left(\begin{matrix}\cos(\pi/N)&-\sin(\pi/N)\\ \sin(\pi/N)&\cos(\pi/N)\end{matrix}\right)\in\mathrm{SO}_{2}(\mathbf{R}).$
@@
Note that $\zeta_{N}\widetilde{F}_{N}$ is another covering map such that $\zeta_{N}\widetilde{F}_{N}(i)=0$. Therefore $\zeta_{N}\widetilde{F}_{N}$ must differ from $\widetilde{F}_{N}$ by a Möbius transformation in the stabilizer of $i$; that is $\zeta_{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\widetilde{r}_{N}\cdot\tau)$ for some $\widetilde{r}_{N}\in\mathrm{SO}_{2}(\mathbf{R})$. We deduce that $\widetilde{F}_{N}(\widetilde{r}^{N}_{N}\cdot\tau)=\zeta_{N}^{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\tau)$, and thus $\widetilde{r}^{N}_{N}\in\mathrm{SO}_{2}(\mathbf{R})$ must also lie in $\widetilde{\Gamma}_{N}$. But $\widetilde{\Gamma}_{N}$ is a free group (due to the fact that $\widetilde{F}_{N}$ is a covering map with no ramification points), and hence $\widetilde{r}^{N}_{N}$ is trivial in $\mathrm{SO}_{2}(\mathbf{R})/\{\pm I\}$, and $\widetilde{r}_{N}$ is a hyperbolic rotation around $i$ of order $N$.
The action of $\mathrm{SO}_{2}(\mathbf{R})$ on $D(0,1)$ under the pullback map is just given by rotation, and hence $\widetilde{r}_{N}$ acts on $D(0,1)$ by a rotation of order $N$. We deduce that $F_{N}(\zeta^{m}z)=\zeta F_{N}(z)$ for some $(m,N)=1$. By taking the derivatives with respect to $q$ of both sides at $q=0$, we have $\zeta^{m}F_{N}^{\prime}(0)=\zeta F_{N}^{\prime}(0)$. Since $F_{N}$ is a covering map, we must also have $F^{\prime}_{N}(0)\neq 0$, and thus $m=1$ and $F_{N}(\zeta z)=\zeta F_{N}(z)$. We thus also deduce (5.1.3) as this action corresponds to the rotation by degree $2\pi/N$ on $D(0,1)$. ∎
@@
To prove this formula, we follow Hempel [Hem88] to get a second order linear ODE whose ratio of two linearly independent solutions gives the (local analytic) inverse of $\widetilde{F}_{N}$ (Lemma 5.1.8). The uniformization maps of Riemann surfaces—and their inverses—do not typically admit explicit solutions in terms of standard functions, but our particular case of interest turns out to be an exception due to the extra symmetries of $\mathbf{C}\smallsetminus\mu_{N}$. We use Lemma 5.1.2 to define a function $G_{N}$ closely related to $F_{N}$ (see Definition 5.1.13) and explicitly find two solutions of the associated linear ODE in terms of hypergeometric functions (Lemma 5.1.14). These solutions allow us to compute the explicit conformal radius for $G_{N}$ and deduce the formula (5.1.5) for $F_{N}$.
Our computation here is very similar to the treatment by Goluzin in [Gol69, §III.1], who also gives the explicit formula for the inverse of $G_{N}$ by hypergeometric functions. See the $q=0$ case of equation (17) and the last paragraph on page 86 of loc. cit. Goluzin more generally computes the Riemann map for the $\mathbf{Z}/N\mathbf{Z}$-rotationally symmetric circular $N$-gon with angles $\pi q$, and explains [Gol69, § II.6] how the $q=0$ case (formula (21) on page 86 of loc. cit.) by Schwarz reflections entails a description of $G_{N}$ and $F_{N}$.
@@
The equation (5.1.9) is more transparent in terms of the Schwarzian derivative $\{\tau,\widetilde{F}_{N}\}$ defined and computed in the proof below:
@@
Proof.
First, since $F_{N}$ is the composition of $\widetilde{F}_{N}$ and a Möbius transformation, we only need to prove the similar assertion for $\widetilde{F}_{N}$. We have a companion uniformization map $1/\widetilde{F}_{N}:\mathbf{H}\rightarrow\mathbf{P}^{1}\smallsetminus\{0,\mu_{N}\}$. (The reason for first considering the reciprocal of $\widetilde{F}_{N}$ is that the standard form considered in [Hem88] is for maps from $\mathbf{H}$ to $\mathbf{P}^{1}\smallsetminus S$ where $S$ is a finite set of points which does not contain $\infty$.)
@@
We now derive $\{\tau,1/\widetilde{F}_{N}\}$ and then $\{\tau,\widetilde{F}_{N}\}$ following [Hem88, §3, §6].
Let $p_{k}=\zeta_{N}^{k}=e^{2\pi ik/N}$ for $k=1,\ldots,N$ and let $p_{0}=0$. We deduce from [Hem88, Theorem 3.1] that the Schwarzian $\{\tau,1/\widetilde{F}_{N}\}$ is given by
@@
where $\psi$ denotes the digamma function: the logarithmic derivative of the Gamma function. Note that the above hypergeometric function is multivalued around $x=0$, but all different branches are accounted for by the branches of the logarithm. ∎
@@
Since $F_{N}$ is a covering map of $\mathbf{C}\smallsetminus\mu_{N}$, then local inverse $\psi_{N}$ is naturally defined on $D(0,1)\subset\mathbf{C}\smallsetminus\mu_{N}$. Moreover, since $F_{N}(1)=1$, we have that $\lim_{x\rightarrow 1}\psi_{N}(x)=1$, where $x\in D(0,1)$ approaches $1$. (A priori, we only conclude that $\lim_{x\rightarrow 1}\psi_{N}(x)$ approaches a cusp of $D(0,1)$, and thus that $|\lim_{x\rightarrow 1}\psi_{N}(x)|=1$; this suffices for the rest of the proof. The more precise statement $\lim_{x\rightarrow 1}\psi_{N}(x)=1$ follows from our assumption $\psi_{N}(0)=0$, either by the description of the fundamental domain further down in Lemma 5.2.1, or more directly by the computation of the rest of the proof that follows, which shows that that $\lim_{x\rightarrow 1}\psi_{N}(x)$ is a positive real number.)
@@
If $N=2$, then $\mathbf{C}\smallsetminus\{\pm 1\}$ is itself biregular to $Y(2)=\mathbf{P}^{1}\smallsetminus\{0,1,\infty\}$ and thus one can find a direct description of the uniformization $\mathbf{H}\rightarrow\mathbf{C}\smallsetminus\{\pm 1\}$ sending $i$ to 0 by $2\lambda(\tau)-1$. In this case, the formulas above specialize to the standard identity $q=e^{-\pi K^{\prime}/K}$ where the elliptic periods $K^{\prime}$ and $K$ are directly related to hypergeometric functions. The only other such case of an incidental isomorphism $\mathbf{C}\smallsetminus\mu_{N}\cong Y(N)$ is $N=3$: this is [Hem88, § 6 Example 5].

### v4

@@ 4.6. An enhancement of Ihara’s Lemma
(4.6.1) $H^{1}(\Gamma(N),\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q}),$
@@
    The twisted restriction map coming from viewing $\psi\in H^{1}(\Gamma(N),\mathbf{F}_{q})$ as a map $\Gamma(N)\rightarrow\mathbf{F}_{q}$ and then considering the map
    $A\psi:\Gamma(N)\cap\Gamma_{0}(p)\rightarrow\mathbf{F}_{q},\quad g\mapsto\psi(AgA^{-1}).$
@@
The kernel of the map (4.6.1) lies inside $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})^{2}$.
@@
This version of Ihara’s Lemma was essentially proved by Ribet in [Rib84]. The proof is just an abelian version of Lemma 4.4.1. We recall some of the details. Let $\Phi\subset\Gamma(N)$ be the maximal normal subgroup whose quotient is an elementary $q$-abelian group $T$. Canonically, we have $\Gamma(N)/\Phi\simeq T$ and $H^{1}(\Gamma(N),\mathbf{F}_{q})\simeq\mathrm{Hom}(T,\mathbf{F}_{q})$. The kernel of the map
@@
Exactly as in the proof of Lemma 4.4.1, we deduce from Goursat’s Lemma that the cokernel $\Delta$ arises from two maps from $\Gamma(N)$ to $\Delta$ which agree along $\Gamma(N)\cap\Gamma_{0}(p)$, and thus on their amalgam $\mathrm{SL}_{2}(\mathbf{Z}[1/p])(N)$. Since $\mathrm{SL}_{2}(\mathbf{Z}[1/p])(N)$ has the congruence subgroup property, it thus arises from a congruence quotient of this group at primes away from $p$. But that precisely means that the classes in $\mathrm{Hom}(T,\mathbf{F}_{q})=H^{1}(\Gamma(N),\mathbf{F}_{q})$ become trivial after passing to a congruence subgroup $\Gamma^{\prime}\subset\Gamma(N)$, hence the claim. ∎
@@
also lies inside $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})^{2}$.
@@ 5. The uniformization of $\mathbf{C}\smallsetminus\mu_{N}$
In this section we develop all the particular analytic properties that we need of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ for $N\geq 2$. André has pointed out to us that our two main results here, namely Theorem 5.1.4 and Lemma 5.2.18, appear in work of Kraus and Roth [KR16, Remark 5.1 and Theorems 1.2 and 1.10]. Nevertheless, as our proofs are simplified to cover our current needs, and since the results of Kraus and Roth rely on some previous work of themselves and others, we keep our self-contained exposition as a convenience to the reader, and refer to [ASVV10, KRS11, KR16] and the references there for various further results and a more thorough study of the uniformization of $\mathbf{C}\smallsetminus\mu_{N}$. The reader will also benefit from the material in § III.1 in Goluzin’s book [Gol69], which recovers $F_{N}$ via an explicit computation of the Riemann map of a $\mathbf{Z}/N\mathbf{Z}$-rotationally symmetric circular $N$-gon, taking the case of zero angles and doing Schwarz reflections in the sides of the circular polygon.
@@ 5.1. Schwarzians and the conformal radius
where $\widetilde{\Gamma}_{N}\subset\mathrm{PSL}_{2}(\mathbf{R})$ denotes the Fuchsian group of $\mathbf{C}\smallsetminus\mu_{N}$. The map $\widetilde{F}_{N}$ is unique up to the action of $\mathrm{PSL}_{2}(\mathbf{R})$ by Möbius transformations on the source, which also changes $\widetilde{\Gamma}_{N}$ by conjugation. The cusps of $\widetilde{\Gamma}_{N}$ are the elements $x\in\partial\mathbf{H}=\mathbf{P}^{1}(\mathbf{R})=\mathbf{R}\cup\{i\infty\}$ such that the stabilizer of $x$ under $\widetilde{\Gamma}_{N}$ contains a parabolic element. If $\mathbf{H}^{*}$ denotes the union of $\mathbf{H}$ with the cusps, then $\mathbf{H}^{*}/\widetilde{\Gamma}_{N}$ may be identified with the compactification $\mathbf{P}^{1}$ of $\mathbf{C}\smallsetminus\mu_{N}$. Since $\mathrm{PSL}_{2}(\mathbf{R})$ acts transitively on $\partial\mathbf{H}$, we may assume, after translation by an element of $\mathrm{PSL}_{2}(\mathbf{R})$, that $i\infty$ is a cusp of $\widetilde{\Gamma}_{N}$ and that $\widetilde{F}_{N}(i\infty)=1$. The stabilizer of $i\infty$ in $\mathrm{PSL}_{2}(\mathbf{R})$ consists of Möbius transformations of the form $\tau\rightarrow a\tau+b$ for some $a,b\in\mathbf{R}$. Thus we may pin down $\widetilde{F}_{N}$ and $\widetilde{\Gamma}_{N}$ exactly by further specifying that $\widetilde{F}_{N}(i)=0$.
@@
Note that $F_{N}$ is just the map $\widetilde{F}_{N}$ composed with a conformal isomorphism $D(0,1)\rightarrow\mathbf{H}$ sending 0 to $i$, and hence
@@
(5.1.3) $\widetilde{r}_{N}=\left(\begin{matrix}\cos(\pi/N)&-\sin(\pi/N)\\ \sin(\pi/N)&\cos(\pi/N)\end{matrix}\right)\in\mathrm{PSO}_{2}(\mathbf{R}).$
@@
Note that $\zeta_{N}\widetilde{F}_{N}$ is another covering map such that $\zeta_{N}\widetilde{F}_{N}(i)=0$. Therefore $\zeta_{N}\widetilde{F}_{N}$ must differ from $\widetilde{F}_{N}$ by a Möbius transformation in the stabilizer of $i$; that is $\zeta_{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\widetilde{r}_{N}\cdot\tau)$ for some $\widetilde{r}_{N}\in\mathrm{PSO}_{2}(\mathbf{R})$. We deduce that $\widetilde{F}_{N}(\widetilde{r}^{N}_{N}\cdot\tau)=\zeta_{N}^{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\tau)$, and thus $\widetilde{r}^{N}_{N}\in\mathrm{PSO}_{2}(\mathbf{R})$ must also lie in $\widetilde{\Gamma}_{N}$. But $\widetilde{\Gamma}_{N}$ is a free group (due to the fact that $\widetilde{F}_{N}$ is a covering map with no ramification points), and hence $\widetilde{r}^{N}_{N}$ is trivial in $\mathrm{PSO}_{2}(\mathbf{R})$, and $\widetilde{r}_{N}$ is a hyperbolic rotation around $i$ of order $N$.
The action of $\mathrm{PSO}_{2}(\mathbf{R})$ on $D(0,1)$ under the pullback map is just given by rotation, and hence $\widetilde{r}_{N}$ acts on $D(0,1)$ by a rotation of order $N$. We deduce that $F_{N}(\zeta^{m}z)=\zeta F_{N}(z)$ for some $(m,N)=1$. By taking the derivatives with respect to $q$ of both sides at $q=0$, we have $\zeta^{m}F_{N}^{\prime}(0)=\zeta F_{N}^{\prime}(0)$. Since $F_{N}$ is a covering map, we must also have $F^{\prime}_{N}(0)\neq 0$. We deduce that $\zeta^{m}=\zeta$ and hence also that $F_{N}(\zeta z)=\zeta F_{N}(z)$. We thus also deduce (5.1.3) since it follows that $\widetilde{r}_{N}$ is a (hyperbolic) rotation by $2\pi/N$ degrees around $\tau=i$ in $\mathbf{H}$. ∎
@@
To prove this formula, we follow Hempel [Hem88] to get a second order linear ODE whose ratio of two linearly independent solutions gives the (local analytic) inverse of $\widetilde{F}_{N}$ (Lemma 5.1.8). The uniformization maps of Riemann surfaces—and their inverses—do not typically admit explicit solutions in terms of standard functions, but our particular case of interest turns out to be an exception due to the extra symmetries of $\mathbf{C}\smallsetminus\mu_{N}$. We use Lemma 5.1.2 to define a function $G_{N}$ closely related to $F_{N}$ (see Definition 5.1.14) and explicitly find two solutions of the associated linear ODE in terms of hypergeometric functions (Lemma 5.1.15). These solutions allow us to compute the explicit conformal radius for $G_{N}$ and then derive the corresponding conformal radius for $F_{N}$ given in equation (5.1.5).
Our computation here is very similar to the treatment by Goluzin in [Gol69, § III.1], who also gives the explicit formula for the inverse of $G_{N}$ in terms of hypergeometric functions. See the $q=0$ case of equation (17) and the last paragraph on page 86 of loc. cit. Goluzin more generally computes the Riemann map for the $\mathbf{Z}/N\mathbf{Z}$-rotationally symmetric circular $N$-gon with angles $\pi q$, and explains [Gol69, § II.6] how the $q=0$ case (formula (21) on page 86 of loc. cit.) by Schwarz reflections entails a description of $G_{N}$ and $F_{N}$.
@@
The equation (5.1.9) is more transparent in terms of the Schwarzian derivative:
@@
We recall here the role [dSG16, § IV.1.2] of Schwarz’s departure from infinitesimal projectivity:
$\{w,x\}:=\left(\frac{w^{\prime\prime}}{w^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{w^{\prime\prime}}{w^{\prime}}\right)^{2}=\frac{d^{2}}{dx^{2}}\log{\frac{dw}{dx}}-\frac{1}{2}\left(\frac{d}{dx}\log{\frac{dw}{dx}}\right)^{2},$
the simplest differential operator invariant under all Möbius transformations. It is featured in the ODE [dSG16, Proposition VIII.3.5]
(5.1.10) $\frac{d^{2}y}{dx^{2}}+\frac{1}{2}\{w,x\}y=0$
that can be used to formally represent an unknown function $w$ as the quotient $w=v_{1}/v_{2}$ of the two linearly independent solutions $y=v_{1}:=w/\sqrt{w^{\prime}}$ and $y=v_{2}:=1/\sqrt{w^{\prime}}$ of the second-order linear ODE (5.1.10). Following Poincaré in his ODE approach to the uniformization of Riemann surfaces, we are interested to describe in this way the multivalued holomorphic inverse $w:U\to\mathbf{H}$ to an analytic universal covering map $\mathbf{H}\to U$, in the case that the Riemann surface $U=\mathbf{C}\smallsetminus\{a_{1},\ldots,a_{N}\}$ is the complement of finitely many punctures in the complex plane, and with $x$ taken as some local coordinate of the complex projective line. In this case, by a local analysis near the punctures $\{a_{i}\}\cup\{\infty\}\subset\mathbf{P}^{1}$, the Schwarzian $\{w,x\}\in\mathbf{C}(x)$ is simply a rational function, which is far easier to compute in practice than the map $w$ a priori. The ODE (5.1.10) then furnishes a local analytic description of the requisite inverse map $w=v_{1}/v_{2}$ which can then be analyzed both locally (as in the rest of the current § 5.1) and globally (as in the next § 5.2).
Example 5.1.11 (See also Example 5.1.21).
@@
Proof of Lemma 5.1.8.
First, since $F_{N}$ is the composition of $\widetilde{F}_{N}$ and a Möbius transformation, we only need to prove the similar assertion for $\widetilde{F}_{N}$. By taking reciprocals, we have a companion uniformization map $1/\widetilde{F}_{N}:\mathbf{H}\rightarrow\mathbf{P}^{1}\smallsetminus\{0,\mu_{N}\}$. (The reason for first considering the reciprocal of $\widetilde{F}_{N}$ is that the standard form considered in [Hem88] is for maps from $\mathbf{H}$ to $\mathbf{P}^{1}\smallsetminus S$ where $S$ is a finite set of points which does not contain $\infty$.) This is similar to Example 5.1.11, except now taking $x=\widetilde{F}_{N}$, rather than $x=\lambda$, as the coordinate of the projective line, once again parametrized by $\mathbf{H}$ via our universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$.
By [Hem88, Lemma 3.3], the analytic local inverse map of $1/\widetilde{F}_{N}$ (resp. $\widetilde{F}_{N}$) is, up to a Möbius transformation, the ratio of two linearly independent solutions of the differential equation equation $y^{\prime\prime}+\frac{1}{2}\{\tau,1/\widetilde{F}_{N}\}y=0$ (resp. $y^{\prime\prime}+\frac{1}{2}\{\tau,\widetilde{F}_{N}\}y=0$), where $\{\tau,1/\widetilde{F}_{N}\}$ and $\{\tau,\widetilde{F}_{N}\}$ denote the Schwarzian derivatives. We now compute $\{\tau,1/\widetilde{F}_{N}\}$ and then $\{\tau,\widetilde{F}_{N}\}$ following [Hem88, § 3, § 6]. Let $p_{k}=\zeta_{N}^{k}=e^{2\pi ik/N}$ for $k=1,\ldots,N$ and let $p_{0}=0$. We deduce from [Hem88, Theorem 3.1] that the Schwarzian $\{\tau,1/\widetilde{F}_{N}\}$ is given by
@@
where $\psi$ denotes the digamma function $\psi(x)=d/dx\log\Gamma(x)=\Gamma^{\prime}(x)/\Gamma(x)$. Note that the above hypergeometric function is multivalued around $x=0$, but all different branches are accounted for by the branches of the logarithm.
@@
Since $F_{N}$ is a covering map of $\mathbf{C}\smallsetminus\mu_{N}$, the local inverse $\psi_{N}$ is naturally defined on $D(0,1)\subset\mathbf{C}\smallsetminus\mu_{N}$. Moreover, since $F_{N}(1)=1$, we have that $\lim_{x\rightarrow 1}\psi_{N}(x)=1$, where $x\in D(0,1)$ approaches $1$. (A priori, we only conclude that $\lim_{x\rightarrow 1}\psi_{N}(x)$ approaches a cusp of $D(0,1)$, and thus that $|\lim_{x\rightarrow 1}\psi_{N}(x)|=1$; this suffices for the rest of the proof. The more precise statement $\lim_{x\rightarrow 1}\psi_{N}(x)=1$ follows from our assumption $\psi_{N}(0)=0$, either by the description of the fundamental domain further down in Lemma 5.2.1, or more directly by the computation of the rest of the proof that follows, which shows that $\lim_{x\rightarrow 1}\psi_{N}(x)$ is a positive real number.)
@@
If $N=2$, then $\mathbf{C}\smallsetminus\{\pm 1\}$ is biholomorphic to $Y(2)=\mathbf{P}^{1}\smallsetminus\{0,1,\infty\}$, and a direct description of the uniformization $\mathbf{H}\rightarrow\mathbf{C}\smallsetminus\{\pm 1\}$ sending $i$ to 0 is given by $2\lambda(\tau)-1$. In this case, the formulas above specialize to the standard identity $q=e^{-\pi K^{\prime}/K}$ where the elliptic periods $K^{\prime}$ and $K$ are directly related to hypergeometric functions. The only other such case of an incidental isomorphism $\mathbf{C}\smallsetminus\mu_{N}\cong Y(N)$ is $N=3$: this is [Hem88, § 6 Example 5]. In our notation, these two respective uniformization maps $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ are explicitly F_2 : D(0,1) →C∖{±1},   F_2(z) = 2 λ( i 1-z1+z ) - 1 and
$F_{3}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{3},$
$F_{3}(z)$ $=9\eta\left(9i\frac{(2+\sqrt{3}+3i)z+2+\sqrt{3}-3i}{(-6-\sqrt{3}+3i)z+6+\sqrt{3}+3i}\right)^{3}\eta\left(i\frac{(2+\sqrt{3}+3i)z+2+\sqrt{3}-3i}{(-6-\sqrt{3}+3i)z+6+\sqrt{3}+3i}\right)^{-3}+1,$
where $\eta(z):=q^{1/12}\prod_{n=1}^{\infty}(1-q^{2n})$ with $q:=e^{\pi iz}$ is the Dedekind eta function. These explicit examples confirm our general formula for the derivative:
[…]
$|F_{3}^{\prime}(0)|$ $=\frac{\Gamma(1/6)^{3}}{12\pi^{3/2}}=5810565\ldots >2$$>519842\ldots=\sqrt[3]{16}.$